OpenStudy (anonymous):

Can somebody please prove a+(-a)=0

OpenStudy (anonymous):

A number + it's negation = 0?

OpenStudy (amistre64):



OpenStudy (amistre64):

i believe its an axiom; and axioms need no proof

OpenStudy (jamesj):

Yes, this the definition of what -a means. In the standard axiomatic definition of the real numbers there is an axiom that states For every real number x, there is another real number y such that x + y = 0. And we write y = -x

OpenStudy (anonymous):

Perhaps some further clarification would help. I am trying to prove the following: Define V=span{v1,v2} and S=span{s1,s2}. Suppose W = { x | x = v + s, forall v in V, s in S }. Is W a Vector space? Using the axioms of what a vector space is, if I cannot prove that there exists -x such that x + ( -x ) = 0, how can I prove it is a vector space?

OpenStudy (jamesj):

Ok. But this is easy to prove. First note that V and S are themselves subspaces. Then for any x in W, there is a v in V and and s in S such that x = v + s As V and S are themselves subspaces, =v and -s are members of V and S respectively and the vector y = -v + -s is the additive inverse of x.

OpenStudy (jamesj):

that should be "-v and -s are members ..."

OpenStudy (anonymous):

Thanks mate

OpenStudy (jamesj):

Kein Problem!

OpenStudy (jamesj):

= sure thing